OpenStudy (anonymous):

solve the following equation for x where 0

OpenStudy (mimi_x3):

hint: \(sin2x = 2sinxcosx\)

OpenStudy (anonymous):

yes and

OpenStudy (mimi_x3):

well where are you stuck or show me what you did

OpenStudy (anonymous):

the equation is not solved

OpenStudy (mimi_x3):

\[ 2sinxcosx = cosx => 2sinxcosx - cosx = 0 \] why dont you try and solve it

OpenStudy (anonymous):

take cos(x) you will have \[\Large \cos(x)[2\sin(x)-1]=0\] now you have two equations \[\Large \implies \cos(x)=0\] \[\Large \implies 2\sin(x)-1=0\] solve these two equations.

OpenStudy (anonymous):

if i knew how to solve these equations i wouldn;t be asking for help

OpenStudy (anonymous):

to this stage i know, but how to solve from here?

OpenStudy (mimi_x3):

\[\cos(x) = 0 => cosx = \cos\left(\frac{\pi}{2}\right) => x = \frac{\pi}{2} +n*2\pi , n \epsilon\mathbb{Z} \] \[sinx = \sin\left(\frac{\pi}{6}\right) => x = \left(-1\right)^{n}*\frac{\pi}{6} +n*\pi , n \epsilon\mathbb{Z} \]

OpenStudy (anonymous):

i've been to wolfram alpha to mimi

OpenStudy (mimi_x3):

lol i didnt use wolfram..

OpenStudy (anonymous):

the answer is correct though but i have to show it using the unit circle, or the graph

OpenStudy (anonymous):

i think using the sin wave

OpenStudy (mimi_x3):